OpenStudy (anonymous):

Find the distance between (6,0) and (4,1)

pooja195 (pooja195):

Hi :) Use the distance formula to find the distance between (6,0) and (4,1) \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

ok what's next

pooja195 (pooja195):

\[\huge d=\sqrt{(4-6)^2+(1-0)^2}\] Solve from here :)

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

4-6=-2 1-0=1

pooja195 (pooja195):

Good job :) Another method would be the on @UnkleRhaukus has shown. \[\huge d=\sqrt{(4-6)^2+(1-0)^2}\] \[\huge d=\sqrt{(-2)^2+(1)^2}\] -2^2=? 1^2=?

OpenStudy (anonymous):

Oh I am not a fan of the graphing method it seems difficult :/ -2^2=4 1^2=1

pooja195 (pooja195):

Ok good so far! Almost done ^_^ \[\huge d=\sqrt{4+1}\] 4+1=?

OpenStudy (anonymous):

5

pooja195 (pooja195):

\[\huge d=\sqrt{5}\] I suggest you leave it in this form :) unless your teacher asked for decimal form!

pooja195 (pooja195):

Do you understand?

OpenStudy (anonymous):

Yes Thx you soooo much! Thanks for explaining and not giving the answer btw

pooja195 (pooja195):

yw \(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (unklerhaukus):

The formula comes from pythagorus, where the sides of the triangle are the difference in co-ordinate values, the distance is the hypotenuse.